# Transfer Learning in Robotics: An Upcoming Breakthrough? A Review of Promises and Challenges

### v2

@@ 1 The Rise of Transfer Learning in Robotics
To evolve seamlessly in the real world, robots must feature outstanding cognitive abilities allowing them to perceive their environment, act and react to achieve various goals, and learn continuously from observation and experience, while coping with changes and uncertainty in the world. The transfer learning paradigm for robotics is a promising avenue to avoid learning from scratch by reusing previously-acquired experience in new situations, similarly to humans. The core idea of transfer learning in robotics, illustrated in Figure 1, is simple: The experience of a robot performing one task in an environment is leveraged to improve the learning process of a (related) task in a different context. However, successfully performing transfer learning in robotics requires tackling three fundamental questions, namely identifying: (1) when — under which conditions — it is advantageous to transfer, (2) what — which knowledge or experience — to transfer, (3) how — with which method — to transfer it. Notice that these fundamental questions also apply to transfer learning in machine learning (Yang et al., 2020) and, to some extent, follow the central questions of imitation learning [^1] The five central questions of imitation learning are who, when, what, and how to imitate, as well as what makes a successful imitation (Dautenhahn and Nehaniv, 2002). (Dautenhahn and Nehaniv, 2002).
In the conceptual example of Figure 1, the experience of the two fixed-based manipulators placing a box on a conveyer belt (left) can be transferred to a humanoid robot executing the same task (middle-top). In this case, the three fundamental questions can be answered as follows:
@@
    When: Although the embodiments of the two systems differ, the upper body of the humanoid resembles the dual-arm manipulator system. Moreover, the task and environment remain the same, thus it is reasonable to assume that experience can be transferred.
@@
    What: The bimanual manipulation strategy of the two fixed-base manipulators can be transferred to the humanoid robot. However, more degrees of freedom must be controlled for the latter. Moreover, the transferred strategy must respect additional constraints, e.g., balance must be ensured to avoid falls.
@@
    How: Optimal control or reinforcement learning may be leveraged to transfer the bimanual control strategy of the two fixed-base manipulators to the humanoid robot.
@@
In other cases, experience may be transferred across tasks. For instance, the bimanual manipulation strategy used when placing a box on a conveyer belt (Figure 1, middle-top) may be transferred to a handover task (middle). Such transfer may also be achieved across different embodiments: The bimanual strategy of the two fixed-based manipulators (left) may directly be transferred to a different robot executing a related task, e.g., to the humanoid robot handing over the box (middle). Here, the three fundamental questions are answered as follows:
  1. 1.
    When: The environment and the manipulated object remain the same for both tasks. In addition, either the same robot is used for both tasks (middle-top to middle), or crucial features (e.g., bimanuality) are common between the two robots executing the two tasks (left to middle). Therefore, it is reasonable to assume that commonalities can be transferred from the placing task to the handover task.
  2. 2.
    What: Bimanual manipulation strategies can be transferred across tasks. In the case where the humanoid robot is used for both tasks, the entire actions may be directly transferred while taking special care of adjusting only task-specific elements.
  3. 3.
    How: Optimal control or reinforcement learning may again be leveraged. Here, it is important to notice that the implementation of the initial task (placing a box) greatly influences the methods that can be used to transfer experience to a second task (handover). In particular, the use of handcrafted strategies to achieve the initial task may reduce the benefits of the transfer, despite the similarities shared between both tasks.
Experience may also be transferred across environments. For example, the bimanual manipulation strategies used by the humanoid robot to place a box on the conveyer belt or to handover the box (Figure 1, middle-top and middle) may be reused by a humanoid robot manipulating a box in space (right). In this case, the fundamental questions can be answered as follows:
  1. 1.
    When: The humanoid robots on Earth and in space have similar embodiments. Moreover, boxes manipulated on Earth and in space have similar characteristics. Therefore, it is reasonable to assume that commonalities can be transferred from Earth to space.
  2. 2.
    What: The bimanual manipulation strategies can be transferred across environments.
  3. 3.
    How: Special care must be taken to handle the different physical rules of the two environments. In this example, controllers on Earth must handle the influence of gravity. This is not required by controllers acting in space.
@@
Finally, in some cases, transferring knowledge may not be beneficial or may even impede the robot performance. For instance, we consider transferring experience from the fixed-based bimanual setup placing a box in a conveyer belt to a humanoid robot kicking a ball. In this case, employing transfer learning might not be beneficial. Instead, the large differences between the robots, tasks, and environments might even cause transfer learning to impede the methods employed to fulfill the ball-kicking task. This clearly showcases the crucial importance of tackling the three fundamental questions when, what, and how to transfer for learning in robotics.
@@
In this position paper, we contend that transfer learning in robotics has the potential to revolutionize the robot learning paradigm by enabling robots to leverage past experience in novel contexts. However, key challenges remain to be addressed to fulfill this potential. In particular, the fundamental questions when — under which conditions —, what — which semantic or sensorimotor information —, how — with which method — to transfer must be thoroughly tackled for transfer learning across robots, environments, and tasks. This position paper reviews the successes of the field and identifies relevant research questions and promising directions paving the way forward. Starting from the definition of transfer learning in the machine learning field, we propose a unified definition of transfer learning in robotics and subsequently build a novel taxonomy of transfer learning in robotics based on the key concepts of robot, task, and environment (see Section 2). Then, we recount successful applications of transfer learning in robotics and show how they align with the proposed taxonomy (Section 3). In Section 4, we outline challenges, as well as promising research directions to tackle them, including abstraction levels and universal representations for transfer learning in robotics, interpretability, benchmarks and simulations, transfer learning metrics, and dangers of negative transfer. Last but not least, we call for actions to address the most immediate roadblocks in Section 5. Overall, the contributions of this paper are twofold: (1) We provide a unified view of transfer learning in robotics by comprehensibly defining the notion of transfer learning in robotics and by introducing the first taxonomy of transfer learning in robotics; (2) We provide a review of promises and challenges in the field of transfer learning in robotics, identifying the most significant roadblocks on the way to unraveling the full potential of transfer learning in robotics.
@@ 2.1 Taxonomy of Transfer Learning in Machine Learning
We observe that Definition 2.1 implies the following hierarchical taxonomy of transfer learning settings illustrated in Figure 4.
@@ 2.2 Taxonomy of Transfer Learning in Robotics
    Robot transfer learning: $\{\mathcal{R}_{S},\mathcal{D},\mathcal{T}\}\to\{\mathcal{R}_{T},\mathcal{D},\mathcal{T}\}$. The goal of this setting is to endow a target robot with the ability to perform a given task known by other source robot(s) in the same environment. Note that the source and target robots may have (very) different morphologies, kinematics, and sensor modalities, leading to different capabilities. For example, Figure 6 illustrates a transfer between the humanoid dual-arm robot ARMAR-III (Asfour et al., 2006) and ALMA (Bellicoso et al., 2019), a four-legged robot equipped with a robotic arm. Moreover, the transfer can happen at different levels, e.g., at the level of joint or task-space controllers, or at the planning level. Instances of robot transfer learning are (i) imitation learning (Schaal, 1999), where a teacher human or robot provides demonstrations of a task to a student robot that learns to reproduce the given task in the same environment, and (ii) (goal-directed) motion retargeting (Dariush et al., 2008; Yin et al., 2023), whose goal is to learn a mapping between different kinematic structures.
@@
    Task transfer learning: $\{\mathcal{R},\mathcal{D},\mathcal{T}_{S}\}\to\{\mathcal{R},\mathcal{D},\mathcal{T}_{T}\}$. This setting aims at leveraging the ability of a robot to perform a given task to learn how to execute a different task in the same environment. The underlying assumption is that the source and target tasks are — to some extent — similar, so that experience can be reused between source and target tasks. For instance, the box tossing and cloth flinging tasks of Figure 6 share similar dynamics characteristics: In both cases, the robot must generate high-velocity dynamic actions to successfully execute the task. Therefore, we may expect that experience about box tossing may be reused by the robot when learning to fling a cloth. Challenges of task transfer learning include inferring which part of the source task experience should be transferred and at which level (joint or task space, planning, etc). Notice that generalizing a given task to an unseen context is a special case of task transfer learning (Mandlekar et al., 2020; Li and Figueroa, 2023). In this case, the model is made compatible with different instances of the same task.
@@ 3.1 Environment Transfer
Change of environment conditions Kramberger et al. (2016) or the complexity of the environment where the task is being executed Vosylius and Johns (2023) provide examples of generalization to a declaratively new environment. However, the environment (domain), can be different in other aspects that go beyond just the setting – e.g. contact conditions or other physical conditions might not be the same Muratore et al. (2022). One example of such is transfer from the simulation-to-reality or sim-to-real.
@@
Potentially unjustly, but transfer learning is in robotics often associated exactly with sim-to-real, where typically experience is obtained in one domain — the simulation —, and exploited to accelerate learning in the transferred domain — the real world. Several reviews cover sim-to-real transfer learning in robotics, i. e., (Muratore et al., 2022; Zhao et al., 2020), affirming the notion of a huge body of work in this field. The gist of sim-to-real lies in the notion that collecting the data for modern (deep) learning and other AI algorithms in the real world is too expensive in terms of time and resources to scale up (Muratore et al., 2022). Therefore, the data is collected in simulation, despite the difference between the real and simulated domains. This difference, referred to as the “reality gap” (Collins et al., 2019a), needs to be overcome for real world execution, which is done using transfer learning. Since collecting data in the real world is so time-consuming and expensive, researchers might change the domain to a different simulation, ending up with sim-to-sim methodologies. These are applied to demonstrate the behavior of transfer learning methodologies.
@@
Different practices have been proposed for sim-to-real transfer learning, starting with realistic modelling (Muratore et al., 2022). No matter how accurate, modelling will never be fully cover all the aspects of the real world (Muratore et al., 2022), thus other approaches have emerged. Domain randomization, such as randomization of image backgrounds, of physical parameters of objects and robot actions, or of controller parameters (Höfer et al., 2021), is a common approach. By randomizing over, for example, physical parameters, the approach tries to cover the entire spectrum of these parameters in the hope that this includes the parameters that describe the real world. Even so, one-shot transfer learning is seldom successful (Zhao et al., 2020), and additional learning is required, for example using reinforcement learning (Ada et al., 2022), back-propagation (Chen et al., 2018) or both (Lončarević et al., 2022). If there are significantly fewer learning iterations in the target domain, the process is called few-shot transfer learning (Ghadirzadeh et al., 2021). Given that more information can be available in the simulation, the notion of privileged learning was introduced, where the privileged information is used to train a high performance policy, which in turn trains a proprioceptive-only student policy (Lee et al., 2020). The idea was very successfully demonstrated in quadrupedal locomotion by more than one group (Lee et al., 2020; Kumar et al., 2022), and is general enough to be applied for very different tasks, such as excavator walking (Egli and Hutter, 2022) and even robotized handling of textiles (Longhini et al., 2022).
@@ 3.2 Task Transfer
Transferring of robot walking from one domain to the other can be considered more than just domain transfer, as walking itself can be different for different environments. Thus, a walking policy learned for one terrain might be effective for walking on a completely different kind of terrain. Moreover, walking is not an isolated instance: If the robot can learn to throw accurately at one target, a modulation of the throwing task to aim at a different target can in fact be considered at the least a different instance of the same task, if not a a different task overall. Such transfers from one (or several) task instances to a new one have been utilized in robotics before, and it was often referred to as generalization. In this sense, for example, fast learning from a small set of demonstrations was applied with nonlinear autonomous dynamical system (DS), which have the ability to generalize motions to unseen contexts (Khansari-Zadeh and Billard, 2011). Similarly, a set of dynamical systems in the form of dynamic movement primitives was used to generalize to transfer knowledge from known situations to unknown in positions (Ude et al., 2010) and in torques (Deniša et al., 2016), probabilistic movement primitives (ProMPs) encode complete families of motions (Paraschos et al., 2013), TP-GMMs adapt to changes of predefined local frames (Calinon, 2018), and Mixture Density Networks adapt a learned motion primitive to new targets specified in a different space (Zhou et al., 2020) [^4] Note that demonstrations and reproductions performed with the same robot avoid the need for robot transfer. This includes demonstrations acquired via kinesthetic teaching or teleoperation.. Generalization was even termed inter-task transfer learning (Fernández et al., 2010). Thus, in a broad sense of Definition 2.2, such approaches already propose solutions for $\{\mathcal{R},\mathcal{D},\mathcal{T}_{S}\}\to\{\mathcal{R},\mathcal{D},\mathcal{T}_{T}\}$, although they were not called transfer learning. Complete skill models were learned from a set of executions also with DNNs (Lončarević et al., 2022). The adaptation of the skill model for a new environment is commonly referred to as transfer learning.

### v3

@@ 1 The Rise of Transfer Learning in Robotics
To evolve seamlessly in the real world, robots must feature outstanding cognitive abilities allowing them to perceive their environment, act and react to achieve various goals, and learn continuously from observation and experience, while coping with changes and uncertainty in the world. The transfer learning paradigm for robotics is a promising avenue to avoid learning from scratch by reusing previously-acquired experience in new situations, similar to humans. The core idea of transfer learning in robotics, illustrated in Figure 1, is simple: The experience of a robot performing one task in an environment is leveraged to improve the learning process of a (related) task in a different context, i.e., in a different environment or executed by a different robot.
To identify when transfer learning is needed and/or warranted, one need to identify the similarities and differences between the two situations. Following the concepts proposed in imitation learning (Dautenhahn and Nehaniv, 2002), we distinguish between three aspects that can be deemed similar or different: The tasks, the environments, and the bodies (a.k.a. the robots in our case).
In the conceptual example of Figure 1, the experience of the two fixed-based manipulators placing a box on a conveyer belt (left) can be transferred to a humanoid robot executing the same task (middle-top). Transfer learning is made possible and easier as the two situations have many aspects in common:
@@
    Tasks: Both tasks are similar as both robots are engaged in moving an object with two arms. The notion of similarity with respect to the object may be relative to the robot’s own size and payload. For instance, a pair of industrial arms may be able to support objects up to $20$ kg whereas a humanoid robot may carry only a quarter of this. Yet, the task and strategy to solve it, namely ensuring coordinated movement of both arms and the trajectory to place an object on a conveyor’s belt, remains the same.
@@
    Environments: The two environments are similar, as in both cases, the object is to be placed on a conveyor belt. We can also safely assume that, in both cases, the object is subjected to the same external forces (gravity, friction).
@@
    Robots: In this case, the two robots differ. Yet, their differences can be broken into two distinct parts. Both robots are endowed with two arms of similar structure. Solely the second robot is on legs and hence faces the additional challenge of controlling its balance when placing the object on the conveyor belt. This aspect may be resolved without actual learning, for instance, through constrained-based optimization, controlling for additional constraints at the center of mass (Bouyarmane et al., 2018). Alternatively, reinforcement learning or adaptive control may be used to fine tune gains (Khadivar et al., 2023).
@@
Hence, in this case, transfer learning is mostly conducted across the different robots, while taking advantage of their similar bimanual structure. In other cases, when the two robots and/or environments differ importantly, experience may be transferred across tasks. For instance, the bimanual manipulation strategy used when placing a box on a conveyer belt (Figure 1, middle-top) may be transferred to a handover task (middle). Such transfer may also be achieved across different tasks and different embodiments: The bimanual strategy of the two fixed-based manipulators (left) may directly be transferred to a different robot executing a related task, e.g., to the humanoid robot handing over the box (middle). Experience may also be transferred across environments. For example, the bimanual manipulation strategies used by the humanoid robot to place a box on the conveyer belt or to handover the box (Figure 1, middle-top and middle) may be reused by a humanoid robot manipulating a box in space (right). In this example, the physical rules of the two environments strongly differ due to the influence of gravity on Earth.
@@
Finally, in some cases, transferring knowledge may not be beneficial or may even impede the robot performance. For instance, we consider transferring experience from the fixed-based bimanual setup placing a box in a conveyer belt to a humanoid robot kicking a ball. In this case, employing transfer learning might not be beneficial. Instead, the large differences between the robots, tasks, and environments might even cause transfer learning to impede the methods employed to fulfill the ball-kicking task. It is important to highlight that, even if some particular actions can be transferred, conflicting goals may lead to negative transfer. This clearly showcases the crucial importance of identifying the commonalities and differences between two situations when applying transfer learning in robotics. Moreover, it highlights the importance of transfer learning metrics that measure not only the transfer quality, but also the transfer gap between robots, tasks, and environments.
@@
In this position paper, we contend that transfer learning in robotics has the potential to revolutionize the robot learning paradigm by enabling robots to leverage past experience in novel contexts. However, key challenges remain to be addressed to fulfill this potential. In particular, the fundamental question of identifying the similarities and differences across tasks, environments, and robots in an automatic manner remains. This position paper reviews the successes of the field and identifies relevant research questions and promising directions paving the way forward. Starting from the definition of transfer learning in the machine learning field, we propose a unified definition of transfer learning in robotics and subsequently build a novel taxonomy of transfer learning in robotics based on the key concepts of robot, task, and environment (see Section 2). Then, we recount successful applications of transfer learning in robotics and show how they align with the proposed taxonomy (Section 3). In Section 4, we outline challenges, as well as promising research directions to tackle them, including abstraction levels and universal representations for transfer learning in robotics, interpretability, benchmarks and simulations, transfer learning metrics, and dangers of negative transfer. Last but not least, we call for actions to address the most immediate roadblocks in Section 5. Overall, the contributions of this paper are twofold: (1) We provide a unified view of transfer learning in robotics by comprehensibly defining the notion of transfer learning in robotics and by introducing the first taxonomy of transfer learning in robotics; (2) We provide a review of promises and challenges in the field of transfer learning in robotics, identifying the most significant roadblocks on the way to unraveling the full potential of transfer learning in robotics.
@@ 2.1 Taxonomy of Transfer Learning in Machine Learning
We observe that Definition 2.1 implies the following hierarchical taxonomy of transfer learning settings illustrated in Figure 4. Note that source space may correspond to a union of multiple sub-source tasks and/or domains.
@@ 2.2 Taxonomy of Transfer Learning in Robotics
    Robot transfer learning: $\{\mathcal{R}_{S},\mathcal{D},\mathcal{T}\}\to\{\mathcal{R}_{T},\mathcal{D},\mathcal{T}\}$. The goal of this setting is to endow a target robot with the ability to perform a given task known by other source robot(s) in the same environment. Note that the source and target robots may have (very) different morphologies, kinematics, and sensor modalities, leading to different capabilities. For example, Figure 6 illustrates a transfer between the humanoid dual-arm robot ARMAR-III (Asfour et al., 2006) and ALMA (Bellicoso et al., 2019), a four-legged robot equipped with a robotic arm. Moreover, the transfer can happen at different actions levels, e.g., at the level of joint or task-space controllers, or at the planning level, and at different perceptual levels, e.g., across different sensors (Tatiya et al., 2020) or sensory modalities (Lee et al., 2020b). Instances of robot transfer learning are (i) imitation learning (Schaal, 1999), where a teacher human or robot provides demonstrations of a task to a student robot that learns to reproduce the given task in the same environment, (ii) (goal-directed) motion retargeting (Dariush et al., 2008; Yin et al., 2023), whose goal is to learn a mapping between different kinematic structures, and (iii) perceptual transfer, where the robots are equipped with different sensory modalities (Silva et al., 2020) such as touch, vision, sound, or olfaction.
@@
    Task transfer learning: $\{\mathcal{R},\mathcal{D},\mathcal{T}_{S}\}\to\{\mathcal{R},\mathcal{D},\mathcal{T}_{T}\}$. This setting aims at leveraging the ability of a robot to perform a given task to learn how to execute a different task in the same environment. The underlying assumption is that the source and target tasks are — to some extent — similar, so that experience can be reused between source and target tasks. For instance, the box tossing and cloth flinging tasks of Figure 6 share similar dynamics characteristics: In both cases, the robot must generate high-velocity dynamic actions to successfully execute the task. Therefore, we may expect that experience about box tossing may be reused by the robot when learning to fling a cloth. Challenges of task transfer learning include inferring which part of the source task experience should be transferred and at which level (joint or task space, planning, etc). Notice that generalizing a given task to an unseen context is a special case of task transfer learning (Mandlekar et al., 2020; Li and Figueroa, 2023). In this case, the model is made compatible with different instances of the same task. Curriculum learning (Narvekar et al., 2020; Shukla et al., 2022) is also a special case of task transfer learning, where a sequence of intermediary tasks of gradually-increasing difficulty is used to learn a complex target task. Finally, task transfer is also particularly considered in the areas of lifelong learning, where task transfer is considered based on a never-ending stream of data, and compositional learning, which focuses on transfer across compositionally-related tasks (Mendez and Eaton, 2023, 2021).
@@ 3.1 Environment Transfer
Change of environment conditions (Kramberger et al., 2016) or the complexity of the environment where the task is being executed (Vosylius and Johns, 2023) provide examples of generalization to a declaratively new environment. However, the environment (domain), can be different in other aspects that go beyond just the setting – e.g. contact conditions or other physical conditions might not be the same (Muratore et al., 2022). One example of such is transfer from the simulation-to-reality or sim-to-real.
@@
Potentially unjustly, but transfer learning in robotics is often associated exactly with sim-to-real, where typically experience is obtained in one domain — the simulation —, and exploited to accelerate learning in the transferred domain — the real world. Several reviews cover sim-to-real transfer learning in robotics, i. e., (Muratore et al., 2022; Zhao et al., 2020), affirming the notion of a huge body of work in this field. The gist of sim-to-real lies in the notion that collecting the data for modern (deep) learning and other AI algorithms in the real world is too expensive in terms of time and resources to scale up (Muratore et al., 2022). Therefore, the data is collected in simulation, despite the difference between the real and simulated domains. This difference, referred to as the “reality gap” (Collins et al., 2019a), needs to be overcome for real world execution, which is done using transfer learning. Since collecting data in the real world is so time-consuming and expensive, researchers might change the domain to a different simulation, ending up with sim-to-sim methodologies. These are applied to demonstrate the behavior of transfer learning methodologies.
@@
Different practices have been proposed for sim-to-real transfer learning, starting with realistic modelling (Muratore et al., 2022). No matter how accurate, modelling will never be fully cover all the aspects of the real world (Muratore et al., 2022), thus other approaches have emerged. Domain randomization, such as randomization of image backgrounds, of physical parameters of objects and robot actions, or of controller parameters (Höfer et al., 2021), is a common approach. By randomizing over, for example, physical parameters, the approach tries to cover the entire spectrum of these parameters in the hope that this includes the parameters that describe the real world. Even so, one-shot transfer learning is seldom successful (Zhao et al., 2020), and additional learning is required, for example using reinforcement learning (Ada et al., 2022), back-propagation (Chen et al., 2018) or both (Lončarević et al., 2022). If there are significantly fewer learning iterations in the target domain, the process is called few-shot transfer learning (Ghadirzadeh et al., 2021). Few-shot transfer learning has notably been applied to sim-to-real transfer in robotics (Bharadhwaj et al., 2019; Shukla et al., 2023). Given that more information can be available in the simulation, the notion of privileged learning was introduced, where the privileged information is used to train a high performance policy, which in turn trains a proprioceptive-only student policy (Lee et al., 2020a). The idea was very successfully demonstrated in quadrupedal locomotion by more than one group (Lee et al., 2020a; Kumar et al., 2022), and is general enough to be applied for very different tasks, such as excavator walking (Egli and Hutter, 2022) and even robotized handling of textiles (Longhini et al., 2022).
@@ 3.2 Task Transfer
Transferring of robot walking from one domain to the other can be considered more than just domain transfer, as walking itself can be different for different environments. For instance, a pacing gait learned to walk on smooth ground might not be stable enough for walking on mountainous terrains. Thus, walking on mountainous terrains can be seen as a novel task, which may benefit from transferring previously-learned gaits adapted to other terrains. Moreover, walking is not an isolated instance: If the robot can learn to throw accurately at one target, a modulation of the throwing task to aim at a different target can in fact be considered at the least a different instance of the same task, if not a different task overall.
Such transfers from one (or several) task instances to a new one have been utilized in robotics before, and were often referred to as generalization. In this sense, for example, fast learning from a small set of demonstrations was applied with nonlinear autonomous dynamical system (DS), which have the ability to generalize motions to unseen contexts (Khansari-Zadeh and Billard, 2011). Similarly, a set of dynamical systems in the form of dynamic movement primitives was used to generalize to transfer knowledge from known situations to unknown in positions (Ude et al., 2010) and in torques (Deniša et al., 2016), probabilistic movement primitives (ProMPs) encode complete families of motions (Paraschos et al., 2013), TP-GMMs adapt to changes of predefined local frames (Calinon, 2018), and Mixture Density Networks adapt a learned motion primitive to new targets specified in a different space (Zhou et al., 2020) [^3] Note that demonstrations and reproductions performed with the same robot avoid the need for robot transfer. This includes demonstrations acquired via kinesthetic teaching or teleoperation.. Generalization was even termed inter-task transfer learning (Fernández et al., 2010).
Task transfer has also been tackled via meta-learning. In the meta-learning setting, a model is trained on a variety of tasks so that new tasks are solved by using none (zero-shot) or only a limited amount (few-shot) of additional training data (Finn et al., 2017; Nichol et al., 2018). For instance, MAML (Finn et al., 2017) was shown to generalize to new goal velocities and directions in the half cheetah and ant locomotion tasks of the Gymnasium benchmark (Towers et al., 2023) faster than conventional approaches. Other meta-learning approaches tackle the transfer problem from a different perspective by learning loss functions (Bechtle et al., 2021, 2022). The meta-learned loss functions generalize to different tasks, thus alleviating the need of designing task-specific losses.
Thus, in a broad sense of Definition 2.2, such approaches already propose solutions for $\{\mathcal{R},\mathcal{D},\mathcal{T}_{S}\}\to\{\mathcal{R},\mathcal{D},\mathcal{T}_{T}\}$, although they were not called transfer learning. Complete skill models were learned from a set of executions also with DNNs (Lončarević et al., 2022). The adaptation of the skill model for a new environment is commonly referred to as transfer learning.